NuPress Valet has a proposed investment with an initial cost of $62 million and cash flows of $12.5 million for 5 years. Debt represents 44 percent of the capital structure. The cost of equity is 13.7 percent, the pretax cost of debt is 8.5 percent, and the tax rate is 34 percent. What is the company’s WACC?

Select one:

a. 10.25%

b. 9.12%

c. 9.75%

d. 10.14%

e. 9.93%
Answer #1

Weight of equity=100-44=56%

After-tax cost of debt=8.5*(1-tax rate)

=8.5*(1-0.34)=5.61%

WACC=Respective costs*Respective weight

=(5.61*0.44)+(13.7*0.56)

=10.14%(Approx).